OpenStudy (anonymous):

Three friends have a nice meal together, and the bill is $25 The three friends pay $10 each, which the waiter gives to the Cashier The Cashier hands back $5 to the Waiter But the Waiter can't split $5 three ways, so he gives the friends one dollar each and keeps 2 dollars as a tip. They all paid $10 and got $1 back. $10-$1 = $9 There were three of them 3 X $9 = $27 If they paid $27 and the waiter kept $2: $27+$2=$29 Where did the other dollar go? $30 - $1 = $29

OpenStudy (anonymous):

I'm having a bit of trouble getting this riddle. Help, please.

OpenStudy (anonymous):

All right. The correct expression does *not* in fact lie directly in the problem, because the latter part does not use the expression. It's not So, we know that they all paid $27, but the waiter received $2 **Of The $27**, so, adding $27 to $2 is non-sensical, since it should actually be a subtraction (I mean, the waiter, I don't think, paid anything).

jimthompson5910 (jim_thompson5910):

Yes that 2 is part of the 27 since 25+5 = 30 25 + 2 + 3 = 30 The 25 is the actual bill, the 2 is part of the amount the three friends paid (9*3 = 27) and the 3 is the change that went back to the three friends (to make them pay $9 each instead of $10 each)